### v1

# Empowering Multi-Robot Cooperation via Sequential World Models

Abstract
Model-based reinforcement learning (MBRL) has shown significant potential in robotics due to its high sample efficiency and planning capability. However, extending MBRL to multi-robot cooperation remains challenging due to the complexity of joint dynamics. To address this, we propose the Sequential World Model (SeqWM), a novel framework that integrates the sequential paradigm into model-based multi-agent reinforcement learning. SeqWM employs independent, sequentially structured agent-wise world models to decompose complex joint dynamics. Latent rollouts and decision-making are performed through sequential communication, where each agent generates its future trajectory and plans its actions based on the predictions of its predecessors. This design enables explicit intention sharing, enhancing cooperative performance, and reduces communication overhead to linear complexity. Results in challenging simulated environments (Bi-DexHands and Multi-Quad) show that SeqWM outperforms existing state-of-the-art model-free and model-based baselines in both overall performance and sample efficiency, while exhibiting advanced cooperative behaviors such as predictive adaptation and role division. Furthermore, SeqWM has been successfully deployed on physical quadruped robots, demonstrating its effectiveness in real-world multi-robot systems. Demos and code are available at: https://github.com/zhaozijie2022/seqwm-marl.
1 Introduction
Model-based reinforcement learning (MBRL) has been widely applied to robotic systems due to its high sample efficiency (Wu et al., 2023) and planning capability (Sun et al., 2023). However, extending MBRL to multi-robot cooperation remains challenging. Early decentralized model-based multi-agent reinforcement learning (MARL) approaches built independent world models for each agent (Egorov & Shpilman, 2022), overlooking inter-agent couplings and hindering coordination. More recent centralized methods, by contrast, assume full observability or unrestricted communication (Chai et al., 2024; Toledo, 2024; Liu et al., 2024b), performing dynamics modeling and policy optimization in the joint space. These methods face challenges related to modeling complexity and communication overhead in robotic systems with high-dimensional observation and action spaces, limiting their deployment in real-world scenarios.
Between centralized and decentralized paradigms, the distributed sequential paradigm has rapidly developed in recent years and demonstrated unique advantages (Khan, 2025). It reformulates multi-agent decision-making as an autoregressive process: agents communicate and act in a certain order, with each updating its policy conditioned on messages and actions from predecessors (Wen et al., 2022; Hu et al., 2025). This design enables more consistent joint reasoning (Ding et al., 2024) and finer-grained credit assignment (Kuba et al., 2022; Wang et al., 2023) without relying on full communication. It also reduces communication overhead from $O(n^{2})$ in centralized settings to $O(n)$ in chain structures, while offering greater robustness under packet loss or disturbances.
Motivated by these advantages, we propose the Sequential World Model (SeqWM), which integrates the sequential paradigm into model-based MARL to structurally decompose dynamics modeling and action planning. For trajectory prediction, the joint dynamics are modeled as sequential agent-wise rollouts along the communication order, where each agent maintains an independent world model and conditions on the predicted trajectories and actions of its predecessors. For action planning, each agent performs multi-step lookahead based on its predecessors’ predictions, thereby preserving cooperative performance while restricting the search to a low-dimensional subspace consistent with the sequential structure.
Figure 1: Behaviors Comparison.
Methods Overhead
MAT (Hu et al., 2025) ndnd
MAZero (Liu et al., 2024b) $n^{2}Hd$
MARIE (Zhang et al., 2025a) $n^{2}KHd$
SeqComm (Ding et al., 2024) $2n^{2}Kd$
SeqWM nHdnHd
Table 1: Communication overhead.222  $n$ is agents number, $d$ is the message dimension, $H$ is prediction horizon, and $K$ is planning iterations.
We evaluate SeqWM in two challenging multi-robot cooperation environments: Bi-DexHands (Chen et al., 2024a) and Multi-Quad (Xiong et al., 2024), and further validate it on physical multi-robot cooperation tasks using two Unitree Go2-W robots. As shown in Figure 1 and Table 2, SeqWM achieves superior cooperative behaviors under comparable or even lower communication overheads. The key highlights of SeqWM are as follows:
  1. (1)
    By integrating the sequential paradigm, SeqWM structurally decomposes joint dynamics and planning, reducing modeling complexity and communication overhead;
  2. (2)
    Through explicit intention sharing, SeqWM enables the emergence of more advanced cooperative behavior patterns;
  3. (3)
    SeqWM outperforms current state-of-the-art methods on standard benchmarks and has been successfully deployed on physical robots.
These results collectively demonstrate that SeqWM, through sequential world modeling and planning, provides an effective pathway for multi-robot cooperation, balancing performance, efficiency, and real-world applicability.
2 Related Work
Model-based RL. In single-agent robotics, model-based RL has shown remarkable success due to its high sample efficiency (Zhou et al., 2024; Lancaster et al., 2024), with approaches such as PlanCP (Sun et al., 2023) and GPC (Qi et al., 2025) leveraging learned dynamics models to predict trajectories and optimize actions for physical robots. In contrast, existing model-based MARL methods often rely on centralized paradigms, hindering their practical deployment in multi-robot systems. For example, MACD (Chai et al., 2024) and CoDreamer (Toledo, 2024) use transformers or GNNs to integrate full state-action across all agents. Recent efforts such as MARIE (Zhang et al., 2025a) explored decentralized dynamics modeling, but still require communication at each prediction step for agent-wise aggregation. Different from these works, SeqWM assigns each agent an independent world model and predicts trajectories sequentially, significantly reducing modeling complexity and communication overhead for real-world deployment.
Sequential Paradigm. Recent studies in MARL have highlighted the advantages of the sequential paradigm (Khan, 2025), which enables fine-grained credit assignment (Kuba et al., 2022), efficient dynamics modeling (Zhang et al., 2025b), and scalable coordination (Xu et al., 2025). For example, MAT (Wen et al., 2022) and PMAT (Hu et al., 2025) model the decision-making process as a sequence prediction problem, employing transformers to autoregressively predict each agent’s actions. A2PO (Wang et al., 2023) and HARL (Zhong et al., 2024; Liu et al., 2024a) further introduce the sequential update scheme that bring clearer interpretability and ensure monotonic improvement. SeqComm (Ding et al., 2024) extends this idea to the communication domain, where agents exchange information in a sequential order, effectively mitigating non-stationarity while reducing communication overhead. Motivated by these benefits, we integrate the sequential paradigm into model-based MARL, leveraging world models to enhance planning and coordination.
3 Preliminaries
Problem Formulation. We model the fully cooperative task as a decentralized partially observable Markov decision process (dec-POMDP), $\mathcal{M}=\langle\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\Omega,\mathcal{P},\mathcal{R},\gamma\rangle$, where $\mathcal{I}=\{v^{1},\dots,v^{n}\}$ is the set of agents, $\mathcal{S}$ is the global state space, $\mathcal{O}=\prod_{i=1}^{n}O^{i}$ is the joint observation space, and $\mathcal{A}=\prod_{i=1}^{n}A^{i}$ is the joint action space. The observation function $\Omega:\mathcal{S}\times\mathcal{I}\rightarrow\mathcal{O}$ defines each agent’s perception of the environment, while the transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ specifies the environment dynamics. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides a shared scalar signal, and $\gamma$ is the discount factor. Each agent $v^{i}$ learns a local policy $\pi^{i}:O^{i}\rightarrow A^{i}$, which maps its observation $o^{i}$ to an action $a^{i}$. The objective is to learn a joint policy $\pi=\prod_{i=1}^{n}\pi^{i}$ that maximizes the expected discounted return $\sum_{\tau=t}^{\infty}\gamma^{\tau}r_{\tau}$.
Sequential Communication and Decision-Making. In many real-world applications, multi-robot systems are often distributed rather than fully decentralized (Negenborn & Maestre, 2014), allowing inter-agent communication to enhance cooperative performance. A Dec-POMDP can thus be extended to a multi-agent POMDP (Oliehoek et al., 2016), where each agent $v^{i}$ receives messages $e^{i}_{t}$ from other agents and updates its policy $\pi^{i}:O^{i}\times E\rightarrow A^{i}$. To balance efficiency and decision quality, agents adopt communication protocols (Yu et al., 2023), defined as $\phi^{i}:O^{i}\times E\times A^{i}\rightarrow E$. Among them, sequential communication is especially popular for its simplicity and effectiveness (Ding et al., 2024). It organizes agents in a certain order, where each agent acts on its own observation and the message from its predecessor, then passes information forward. Formally, the process is defined as:
$a^{i}_{t}=\pi^{i}(o^{i}_{t},e^{i}_{t}),\quad e^{i+1}_{t}=\phi^{i}(e^{i}_{t},o^{i}_{t},a^{i}_{t}),$ (1)
Such a sequential structure naturally motivates us to design a world model that predicts trajectories in the same manner, enabling efficient multi-agent planning.
4 Methodology
In this section, we propose SeqWM, a Sequential World Model for multi-robot cooperation. Unlike existing centralized or decentralized multi-agent world models, SeqWM decomposes the joint dynamics into agent-wise models arranged in a sequence. This design substantially reduces modeling complexity and communication overhead, enabling deployment in real-world scenarios.
4.1 Sequential Multi-Agent World Model
Figure 2: Comparison of SeqWM’s distributed sequential paradigm with existing centralized/decentralized paradigms.
Decomposed Dynamics Modeling. At each timestep $t$, the observation-action pair of a single agent $(o^{i}_{t},a^{i}_{t})$ can be regarded as a token, and the entire multi-agent system as a sequence of such tokens. This perspective reformulates multi-agent joint dynamics as a sequence modeling problem: given the token sequence $(o^{1}_{t},a^{1}_{t}),\dots,(o^{n}_{t},a^{n}_{t})]$, the model generates the next-step outcomes $(o^{1}_{t+1},r^{1}_{t+1}),\dots,(o^{n}_{t+1},r^{n}_{t+1})]$. Unlike existing centralized world models (Chai et al., 2024; Zhang et al., 2025a), which assume full communication and fuse all tokens simultaneously for prediction, as shown in Figure 2, our method adopts an autoregressive paradigm. In this setup, agent $v^{1}$ first predicts $(o^{1}_{t+1},r^{1}_{t+1})$ from its local information $(o^{1}_{t},a^{1}_{t})$, and passes the result to $v^{2}$. Subsequently, each agent $v^{i}$ conditions on its own observation–action pair $(o^{i}_{t},a^{i}_{t})$ together with the predictions of all predecessors $\{(o^{j}_{t+1},r^{j}_{t+1})\}_{j<i}$ to produce $(o^{i}_{t+1},r^{i}_{t+1})$. Such a sequential design reduces communication frequency and modeling complexity in a structured, scalable manner, making the approach well-suited for real-world deployment.
Multi-Agent World Model. As noted in Introduction, multi-robot cooperative tasks involve high-dimensional observation and action spaces, making it unsuitable to use reconstructing raw observations as the learning objective of the world model. Following (Hafner et al., 2025), we remove the explicit decoder and instead perform dynamics prediction entirely in a latent space. To facilitate distributed deployment, each agent maintains an independent world model without parameter sharing. Let $z^{i}_{t}$ denote the latent state of agent $v^{i}$ at timestep $t$, the world model can be defined as follows:
$\begin{array}[]{rrl}\textrm{Encoder:}&z^{i}_{t}&=E^{i}\left(o^{i}_{t}\right)\\ \textrm{Dynamics:}&\hat{z}^{i}_{t+1}&=D^{i}\left(z^{i}_{t},a^{i}_{t},e^{i}_{t}\right)\\ \textrm{Reward:}&\hat{r}_{t+1}&=R^{i}\left(z^{i}_{t},a^{i}_{t},e^{i}_{t}\right)\\ \textrm{Communication:}&e^{i+1}_{t}&=e^{i}_{t}\oplus\left(z^{i}_{t},a^{i}_{t}\right)\\ \textrm{Critic:}&\hat{q}^{i}_{t}&=Q^{i}\left(z^{i}_{t},a^{i}_{t},e^{i}_{t}\right)\\ \textrm{Actor:}&\hat{a}^{i}_{t}&=\pi^{i,\textrm{Act}}\left(z^{i}_{t},e^{i}_{t}\right).\end{array}$ (2)
All modules in SeqWM are implemented using MLPs, ensuring architectural simplicity and consistency. For communication function, we adopt concatenation operator $\oplus$ to facilitate modular training.333 We implement the concatenation using a mask-based scheme to ensure consistent dimensionality, with details provided in Appendix A.1. As shown in Section 5.4, this concise design achieves more stable training performance compared to alternatives such as cross-attention and recurrent neural networks (RNNs).
Learning Objective. Let $\theta_{E}$, $\theta_{D}$, $\theta_{R}$, $\theta_{Q}$, and $\psi$ denote the parameters of the encoder, dynamics predictor, reward predictor, critic, and actor, respectively. Following the self-supervised training framework (Hafner et al., 2025; Hansen et al., 2024), the loss functions can be defined as:
$\mathcal{L}^{i}(\theta)=\sum_{t}^{H}\lambda^{t}\left(\underbrace{\left\|\hat{z}^{i}_{t+1}-\textrm{sg}(z^{i}_{t+1})\right\|^{2}}_{{\color[rgb]{0.83984375,0.15234375,0.15625}\textrm{dynamics loss for}~\theta_{D},\theta_{E}}}+\underbrace{\textrm{Soft-CE}\left(\hat{r}^{i}_{t},r_{t}\right)}_{{\color[rgb]{0.83984375,0.15234375,0.15625}\textrm{reward loss for}~\theta_{R},\theta_{E}}}+\underbrace{\textrm{Soft-CE}\left(\hat{q}^{i}_{t},G_{t}\right)}_{{\color[rgb]{0.83984375,0.15234375,0.15625}\textrm{Q Loss for}~\theta_{Q},\theta_{E}}}\right),$ (3)
where $\theta=\{\theta_{E},\theta_{D},\theta_{R},\theta_{Q}\}$, $H$ is the prediction horizon, $\lambda\in(0,1]$ is a constant that balances the contribution of each rollout step, $r_{t}$ is the ground-truth reward, $G_{t}$ is the TD target, and $\hat{z}^{i}_{t+1}=D^{i}\left(z^{i}_{t},a^{i}_{t},e^{i}_{t}\right)$ is the predicted latent state. The loss can be backpropagated to the encoder via $z^{i}_{t}=E^{i}\left(o^{i}_{t}\right)$, so do the dynamics and reward. The latent target $z^{i}_{t+1}=E^{i}\left(o^{i}_{t+1}\right)$ is detached with the stop-gradient operator $\textrm{sg}(\cdot)$ to prevent cyclic gradient flow and gradient degradation. Soft Cross-Entropy loss is used to match the discretized reward and Q-value predictions, with details provided in Appendix B.3. Additionally, to ensure modularity and scalability, each agent’s world model is trained independently, and the loss can not be backpropagated through the communication channel.
Based on Eq. (3), the encoder learns a compact latent space, while the dynamics and reward predictors minimize prediction errors in this space, ensuring alignment with real environment dynamics. The actor generates initial action estimates in the latent space to warm-start planning and is trained using the Heterogeneous-Agent Soft Actor-Critic (HASAC) (Haarnoja et al., 2018; Liu et al., 2024a) algorithm:
[…]
where $\hat{a}^{i}_{t}=\pi^{i,\textrm{Act}}\left(z^{i}_{t},e^{i}_{t}\right)$, $\alpha$ is the entropy coefficient, and $\mathcal{H}[\cdot]$ denotes the entropy function.
Sequential Update Scheme. We adopt the sequential update scheme (Kuba et al., 2022; Zhong et al., 2024) to train world models in a manner aligned with its autoregressive structure. When training the agent $v^{i+1}$, its inputs are conditioned on the predictions of the first $i$ agents, produced by their most recently updated models. This preserves the sequential dependency, ensuring predictions exploit the most up-to-date outputs, which stabilizes training and improves monotonicity across agent indices.
Random Masking. Inspired by Masked AutoEncoders (He et al., 2022) and their applications in MARL (Kang et al., 2025), we randomly permute the communication order among agents and allow each agent to skip communication with a certain probability. This random masking simulates realistic communication interruptions and forces the world model to robustly adapt to communication uncertainties, significantly enhancing the model’s resilience against communication failures.
4.2 Planning with Sequential Communication
Although Eq. (2) includes an actor, it does not serve directly as an explicit decision policy; instead, it provides initial action estimates for the planner. We next propose a sequential multi-agent planner based on Model Predictive Path Integral (MPPI) (Williams et al., 2015) that leverages the predictions of world models to optimize each agent’s action. In this framework, the actor contributes by narrowing the action search space to promising regions, while the planner ensures robust long-term decision-making and corrects suboptimal proposals from the actor. 444 Appendix C.3 provides an ablation analysis illustrating the complementary roles of the actor and planner.
Figure 3: Sequential multi-agent planner: each agent optimizes its action sequence via local world model rollouts and critic evaluations, then passes the optimized trajectory to the next agent for efficient sequential cooperation.
At each timestep $t$, agent $v^{i}$ samples $N$ candidate action sequences of horizon $H$, denoted $a^{i}_{t:t+H}$, from the initial distribution guided by the actor. Conditioned on its latent state and the received message, the agent performs latent rollouts with its local world model to predict future trajectories. The value of each trajectory is then estimated as
$V^{i}_{t+H}=\gamma^{H}Q^{i}(\hat{z}^{i}_{t+H},a^{i}_{t+H},e^{i}_{t+H})$ (5)
$+\sum_{h=t}^{t+H-1}\gamma^{h-t}R^{i}(\hat{z}^{i}_{h},a^{i}_{h},e^{i}_{h}),$
where $V^{i}_{t+H}$ represents the value estimate of a sampled action sequence, computed as the sum of predicted rewards over the horizon plus the terminal value given by the critic. Then, candidate sequences are ranked according to their evaluated values, and the highest-scoring subset is selected as the elite set. The action distribution is updated toward the statistics of these elite trajectories, thereby concentrating future sampling around high-value regions and progressively refining the search space. 555 The details of the planner and its implementationare provided in Appendix A.1 and  B.
After several iterations until convergence, the optimized action sequence and predicted trajectory are transmitted as a communication message to the next agent, which repeats the same planning procedure. This sequential communication–planning paradigm substantially reduces communication overhead and enhances multi-agent cooperation efficiency through explicit intention sharing.
Low-pass Action Smoothing. To prevent mechanical wear caused by abrupt action changes during real-world execution, we integrate a low-pass filtering strategy (Kicki, 2025) into the planner. In each planning iteration, sampled action sequences are filtered along the temporal dimension to suppress high-frequency fluctuations. This smoothing enforces gradual action transitions across timesteps, reducing control discontinuities and promoting stable, consistent behavior on physical robots. Further details and formulations are provided in Appendix A.2.
Heuristic Early-Stopping. Considering the computational constraints of real robotic platforms, we design a motion-planning heuristic that terminates iterations when the Kullback-Leibler (KL) (Kullback & Leibler, 1951) divergence between consecutive action distributions falls below a threshold. This early-stopping criterion mitigates diminishing returns (Kobilarov, 2012), reducing computation while preserving plan quality. Further details and experiments regarding this design are provided in Appendix C.4.
Communication Cache. Inspired by Q-chunking (Li et al., 2025) which reuses temporally extended action units to improve decision efficiency, we introduce a communication cache that stores the predicted messages from the previous agent, enabling the current agent to retrieve them when communication fails. For instance, if communication fails at $t+1$, agent $v^{i+1}$ retrieves the cached message $z^{i}_{t+1}=D^{i}(E^{i}(o^{i}_{t}))$ from agent $v^{i}$ instead of the ideally updated message $\hat{z}^{i}_{t+1}=E^{i}(o^{i}_{t+1})$.
5 Experiments
Environments. We evaluate SeqWM and baselines in two challenging multi-robot cooperative environments: Bimanual Dexterous Hands (Bi-DexHands) (Chen et al., 2024a) and Multi-Quadruped Environment (Multi-Quad) (Xiong et al., 2024). In Bi-DexHands, two agents control a pair of dexterous hands to accomplish high-dimensional manipulation tasks (up to $\mathcal{O}\in\mathbb{R}^{229},\mathcal{A}\in\mathbb{R}^{26}$). In Multi-Quad, multiple quadruped robots collaborate to solve coordination tasks, and we further deploy SeqWM on real Unitree Go2-W robots to assess its sim-to-real transfer.
5.1 Comparisons
Baselines. We select several competitive baselines, including: HASAC (Liu et al., 2024a), a state-of-the-art model-free method extending SAC to multi-agent settings; MARIE (Zhang et al., 2025a), a model-based method employing a Transformer for dynamics prediction; MAT (Wen et al., 2022), a method adopting the sequential decision-making paradigm; and MAPPO (Yu et al., 2022), a most widely used algorithm, included as a general-purpose baseline.
Figure 4: Performance comparisons of SeqWM with other baselines. Task in Bi-DexHands report the episode return, while Multi-Quad (gray background) reports success rate. Bold lines indicate the mean over multiple seeds, with shaded regions denoting the 95% confidence intervals.
Results on Bi-DexHands. We select six representative tasks, including object transfer tasks (Over,CatchAbreast,CatchOver2Underarm), which require the two hands to transfer an object under different relative positions and grasping postures; and functional manipulation tasks (BottleCap,Pen,Scissors), which involve precise bimanual operations to achieve specific functional goals, such as opening a bottle cap, removing a pen lid, or spreading a pair of scissors.
As shown in Figure 4, SeqWM achieves higher asymptotic returns and faster convergence across all tasks. In several tasks (Over, CatchOver2Underarm, Scissors), SeqWM reaches near-optimal performance within 2–4M steps, while baselines require far more interactions or fail to match it. In more challenging tasks (Pen, CatchAbreast), SeqWM steadily improves and achieves the highest final returns with lower variance, demonstrating stability. Remarkably, in BottleCap, SeqWM is the only method that successfully grasps the bottle body and opens the cap.
Results on Multi-Quad. The Multi-Quad tasks require two quadruped robots to cooperate under spatial constraints and dynamic interactions. In Gate, the robots are required to pass through a narrow gate as quickly as possible without collision 666 We extend Gate to 3–5 agents in Appendix C.1 to evaluate SeqWM’s scalability.. In PushBox, they jointly push a large box to a designated target location. In Shepherd,, the two quadruped robots (as sheepdogs) cooperatively guide another robot (as sheep) to a target area (as sheep pen).
In Gate and Shepherd, it rapidly approaches near-100% success rates within the early phase, significantly surpassing baselines in terms of sample efficiency. In PushBox, SeqWM is the only method capable of completing the task, whereas all baselines fail to push the box to the target. This superior performance stems from SeqWM’s sequential world model, which enables each agent to plan actions conditioned on its predecessors’ intentions, thereby enhancing coordination.
5.2 Empowered Cooperative Behaviors
We further visualize the behaviors learned by SeqWM, showing that it not only learns stable policies in high-dimensional state and action spaces, but also achieves advanced cooperative behaviors, including predictive adaptation, role division, temporal alignment, and intention sharing.
Bi-DexHands Behaviors. Trajectory visualizations in Figure 5 show that SeqWM acquires sophisticated coordination strategies. In Catch-Over2Underarm, the receiving hand anticipates the object’s motion based on predictive communication from its partner and pre-adapts its grasp posture, enabling a precise and stable catch near the target. In Pen, both hands achieve near-perfect temporal alignment, grasping the pen body and cap almost simultaneously to complete the removal action efficiently. These behaviors stem from SeqWM’s combination of world model–based planning and sequential decision-making. The leading agent first predicts and plans its actions, explicitly transmitting future trajectories to its partner. This intention sharing enables the receiving agent to act on foresight rather than delayed perception, producing synchronized, task-specific coordination and enhancing collaborative efficiency.
Figure 5: Trajectory visualizations of the tasks selected for Bi-DexHands with SeqWM.
Multi-Quad Behaviors. Figure 6 illustrates the coordinated behavior learned by SeqWM in PushBox. From $t=1$ (frame A) to $t=2$ (frame B), the quadrupeds navigate to opposite sides of the box, establishing an effective pushing configuration. Between $t=2$ (frame B) and $t=4$ (frame D), both maintain high positive $x$-axis velocities, indicating sustained forward pushing force. At $t=3$ (frame C), Robot 2 produces a downward $y$-axis velocity, adjusting the push direction toward the target, while Robot 1 gradually increases its negative $y$-axis velocity to assist with directional control. As the box nears the target, Robot 1 reduces its $x$-axis velocity to avoid overshooting. These behaviors demonstrate that SeqWM supports not only effective force coordination for object propulsion but also fine-grained directional adjustments, resulting in precise and efficient task completion.
Figure 6: Behavior visualizations in PushBox. The first row shows the execution process, where the box is significantly larger than the robots, requiring coordinated efforts from both quadrupeds to complete the task. The second row visualizes the trajectories of the robots and the box, with the right side showing the x-axis and y-axis velocities and orientations of each robot.
5.3 Real-World Deployment
Real-World Results. The real-world experimental setup is detailed in Appendix C.5, and the results are shown in Figure 7.
Figure 7: Real-world results of multi-robot cooperation experiments. The trajectories of Robot 1, Robot 2, and the Sheep are marked in different colors.
In PushBox, the two quadrupeds approach the box from opposite sides and coordinate their pushing forces and directions to move it toward the target. Between frames (D-F), Robot 1 moves forward to provide the main pushing force, while Robot 2 makes slight lateral adjustments to steer the box. The overall pushing pattern, including the division of roles and the gradual directional adjustments, closely matches the behavior observed in simulation, confirming a successful sim-to-real transfer.
In Gate, two clear yielding events are observed. Between frames (C-D), Robot 1 slows down and waits for Robot 2 to pass first, demonstrating priority management in constrained spaces. After crossing (frames E-F), Robot 2 actively veers aside to leave sufficient space for Robot 1, enabling smooth passage without collisions. These behaviors reflect SeqWM’s trajectory prediction and intention-sharing capabilities, allowing natural, efficient yielding without high-frequency communication.
In Shepherd, Robot 1 accelerates between frames (A-B), causing the Sheep to move left. To prevent the Sheep from hitting the left gate frame, Robot 1 retreats while Robot 2 advances between frames (C-D). This maneuver drives the Sheep away from Robot 2 and into the target area. The sequence highlights SeqWM’s capacity for predictive coordination and adaptive role allocation, where the one agent’s motion influences the sheep robot’s response and the another agent adapts accordingly to achieve the common goal.
5.4 Ablation Studies
Sequential Sample Generation. To evaluate the contribution of the sequential world model in SeqWM, we replace it with centralized and decentralized architectures, ensuring all models have an equal number of parameters for a fair comparison. Using BottleCap, we collect 50K environment steps with random actions and train each model for 2.5K steps using the loss in Eq. (3). After training, we gather 1K additional steps to measure dynamics and reward prediction errors across different horizons. As shown in Figure 8, the sequential and centralized models achieve similarly low errors, both substantially outperforming the decentralized model. The results confirm the advantage of sequential prediction, where each agent conditions its output on the predictions of its predecessors, yielding more accurate and coherent rollouts.
Figure 8: Accuracy of the world model prediction.
Figure 9: Ablation study of the communication function in SeqWM.
Communication Function. We replaced the concat communication function in SeqWM with alternative fusion mechanisms, including MLP, cross-attn, and RNN, and evaluated them on the BottleCap task. The results in Figure 9 show that the simplest concat approach achieves the highest and most stable performance. This advantage stems from two factors: (i) concat preserves the complete communication content, allowing the dynamics and reward predictors to autonomously identify and exploit the most informative features during training; and (ii) it introduces no additional learnable parameters, thereby maintaining stable gradient propagation in long-horizon prediction. Moreover, we observe that RNN-based fusion even underperforms the no-communication baseline (dec), which we attribute to its sensitivity to input ordering—an undesirable property in multi-agent communication scenarios lacking a fixed semantic sequence.
5.5 Additional Experiments
To further validate the effectiveness of SeqWM, we present the following additional analyses:
  1. 1)
    Visualizations the learned behaviors on all other tasks in Appendix C.2;
  2. 2)
    An ablation analysis of the actor to illustrate its complementary role with the planner in Appendix C.3;
  3. 3)
    An analysis of the planner’s time cost and early-stopping heuristic in Appendix C.4;
  4. 4)
    An evaluation of SeqWM’s scalability with respect to the number of agents in Appendix C.1.
6 Conclusion
This paper presented SeqWM, a novel model-based MARL framework that integrates the sequential paradigm into world model learning and planning. By structurally decomposing joint dynamics into autoregressive, agent-wise models, SeqWM offers a principled approach that reduces modeling complexity, lowers communication overhead, and naturally enables intention sharing through predicted trajectories. This methodological innovation not only improves scalability but also facilitates the emergence of advanced cooperative behaviors such as predictive adaptation, temporal alignment, and role division. Extensive experiments in Bi-DexHands and Multi-Quad show that SeqWM achieves state-of-the-art performance with superior sample efficiency, while real-world deployment on quadruped robots confirms that these cooperative behaviors transfer reliably from simulation to physical platforms. Beyond empirical results, SeqWM demonstrates that sequential paradigms provide a scalable and theoretically grounded principle for structuring multi-agent cooperation, paving the way for more robust and efficient deployment of cooperative policies in real-world systems.
Appendices
[…]
$\printcontents[appendices]$l0
Appendix A Multi-Agent Planner
Figure 10: The multi-agent planner in SeqWM.
A.1 Planning Process
As shown in Figure 10, at timestep $t$, the action planning process for agent $v^{i}$ can be divided into the following steps:
S1 - Communication. Agents are organized to exchange messages in a sequential manner. Specifically, agent $v^{i}$ receives a message $e^{i}_{t}$ that aggregates the predicted latent states and planned actions from all its predecessors:
$e^{i}_{t}=\begin{cases}\emptyset,&i=1,\\ \bigoplus_{j<i}\left(\hat{z}^{j}_{t},a^{j}_{t}\right),&i>1,\end{cases}$ (6)
where $\oplus$ denotes concatenation. To implement this efficiently, we employ a masking-based concatenation scheme: a fixed-length vector of dimension $n\times(|\mathcal{A}|+d_{z})$ is pre-allocated, where $n$ is the number of agents, $|\mathcal{A}|$ and $d_{z}$ are the action and latent dimensions. Agent $v^{1}$ maintains an empty message, while subsequent agents sequentially fill in their designated slots with their own predictions $(\hat{z}^{i}_{t},a^{i}_{t})$ in addition to forwarding the received content. This design ensures that information is progressively accumulated along the communication chain with linear complexity in the number of agents.
S2 - Action Sampling. The planner samples $N$ candidate action sequences from two sources. We sample $N_{p}$ candidate action sequences from a diagonal Gaussian distribution $a^{i}_{t:t+H}\sim\mathcal{N}\left(\mu^{i}_{t:t+H},(\sigma^{i}_{t:t+H})^{2}I\right)$, where $\mu^{i}_{t:t+H},\sigma^{i}_{t:t+H}\in\mathbb{R}^{|\mathcal{A}|\times H}$ represent the mean and standard deviation of the $H$-step horizon actions. Additionally, we sample $N_{a}$ action sequences directly from the actor module $\hat{a}^{i}_{h}\sim\pi^{i,\textrm{Act}}(\cdot|o^{i}_{h},e^{i}_{h}),h=t:t+H$, and combine these two sets of action sequences to form $N$ candidate action sequences.
S3 - World Model Prediction. Following sampling, the world model predicts $H$-step trajectories for each sampled action sequence using Eq. (2), generating $N$ predicted sequences $\Gamma=\{(\hat{z}^{i}_{h},a^{i}_{h},\hat{r}^{i}_{h})\}_{h=t:t+H}$.
S4 - Value Evaluation. Each predicted trajectory is assigned a value via the $H$-step return, combining the short-term cumulative predicted reward with the terminal value from the critic:
$V^{i}_{\Gamma}=\gamma^{H}Q^{i}(\hat{z}^{i}_{t+H},a^{i}_{t+H},e^{i}_{t+H})+\sum_{h=t}^{t+H-1}\gamma^{h-t}\hat{r}^{i}_{h}.$ (7)
S5 - Action Optimization. The candidate action sequences are ranked by their evaluated values, and the top $M$ are chosen as the elite set $\Gamma^{*}$. The parameters of the action distribution are updated based on the elite set using:
$\mu^{i,(k+1)}_{t:t+H}$ $=\frac{\sum_{m=1}^{M}\alpha_{m}\Gamma_{m}^{*}}{\sum_{m=1}^{M}\alpha_{m}},\ \ \ \ \sigma^{i,(k+1)}_{t:t+H}$ $=\sqrt{\frac{\sum_{m=1}^{M}\alpha_{m}\left(\Gamma_{m}^{*}-\mu^{i,(k+1)}_{t:t+H}\right)^{2}}{\sum_{m=1}^{M}\alpha_{m}}},$ (8)
where the weights are generated based on the evaluated values as $\alpha_{m}=\exp\left[\tau\left(V_{\Gamma_{m}^{*}}-\max_{m\in M}V_{\Gamma_{m}^{*}}\right)\right]$, with $\tau$ being the temperature coefficient.
Iteration. For the default setting, the above process are iterated $K$ times to derive the final action distribution. If the early-stopping heuristic is applied, after each iteration, we check whether the action optimization has converged by evaluating the KL divergence between the current and previous action distributions, $\mathbb{D}_{KL}(\mathcal{N}^{(k+1)}\|\mathcal{N}^{(k)})<\eta$, where $\eta$ is a small threshold.
The detailed hyperparameters used in the model-based planner are summarized in Table 3.
A.2 Low-Pass Action Smoothing
In real-world robotics, high-frequency changes in control inputs can cause severe mechanical impacts, accelerating wear and reducing execution stability. Therefore, many studies in reinforcement learning and motion planning incorporate action-smoothing constraints, such as adding penalties on differences between consecutive actions during policy updates (Aractingi et al., 2023; Christmann et al., 2024; Wang et al., 2025), introducing regularization in policy networks (Chen et al., 2024b; Song et al., 2025), or filtering noise in trajectory optimization (Pinneri et al., 2021; Vlahov et al., 2024; Kicki, 2025), to reduce jitter and improve executability. Inspired by these methods, we apply frequency-domain low-pass filtering directly to the sampled action noise in our planner, explicitly suppressing the high-frequency components of the actions.
Figure 11: Visualization of the low-pass filter. (Left): Amplitude–frequency response of the low-pass filter. (Middle): Filtering effects on random signals at 20% cutoff ratios, the different colors represent different action dimensions. (Right): Effects of low-pass filtering on control commands in the PushBox task, with different colors representing different agents.
Specifically, during action sampling, we first sample noise from a standard normal distribution, apply low-pass filtering, and then add the filtered noise to the action mean to generate candidate action sequences. We use a Butterworth filter with $o_{\mathrm{LBF}}=1$, whose transfer function and amplitude–frequency response are given by:
$H(s)=\frac{2\pi f_{c}}{s+2\pi f_{c}},\quad|H(f)|=\frac{f_{c}}{\sqrt{f^{2}+f_{c}^{2}}},$ (9)
where $f_{c}$ is the low-pass cutoff frequency. As show in Figure 11, the amplitude–frequency response shows that the high-frequency components are exponentially attenuated (approaching linear decay in logarithmic coordinates). The corresponding discrete-time difference equation, obtained via bilinear transformation, is
$y[t]=\frac{1-\beta}{2}(x[t]+x[t-1])-\beta y[t-1],\quad\beta=\frac{1-\tan\left(\nicefrac{{\pi f_{c}}}{{f_{s}}}\right)}{1+\tan\left(\nicefrac{{\pi f_{c}}}{{f_{s}}}\right)},$ (10)
where $f_{s}$ is the sampling frequency, i.e., the frequency of the control signal.
Appendix B Implementation Details
B.1 Pseudocode
Algorithm 1 Model Training
Input: replay buffer $\mathcal{B}$, parameterized networks $\theta_{E},\theta_{D},\theta_{R},\theta_{Q},$ and $\psi$ for encoder, dynamics predictor, reward predictor, critic, and actor, respectively;
for episode $1,2,3,\dots,$ do
  for step $t=1,2,3,\dots$ do
   Get real data $([o^{i}_{t}]_{i=1:n},[a^{i}_{t}]_{i=1:n},r_{t},[o^{i}_{t+1}]_{i=1:n})$ by interacting with the environment
   Add transition into buffer: $\mathcal{B}=\mathcal{B}\cup([o^{i}_{t}]_{i=1:n},[a^{i}_{t}]_{i=1:n},r_{t},[o^{i}_{t+1}]_{i=1:n})$
  end for
  for epoch $1,2,3,\dots,$ do
   Sample trajectories from $\mathcal{B}$
   Update $\theta_{E},\theta_{D},\theta_{R},\theta_{Q}$ by minimizing Eq. (3)
   Update $\psi$ by minimizing Eq. (4).
  end for
end for
Algorithm 2 Model Planning
Input: learned parameters $\theta_{E},\theta_{D},\theta_{R},\theta_{Q},\psi$, hyperparameters $H,K,\tau,N_{p},M,N_{a}$, initial distribution;
for step $t=1,2,3,\dots$ do
  for agent $i=1,2,\dots,n$ do
   Get environment observation $o^{i}_{t}$ and encode it to latent space: $z^{i}_{t}=E^{i}(o^{i}_{t})$
   if $i>1$ then
     Retrieve the message from the previous agent $e^{i}_{t}=\bigoplus_{j<i}\left(\hat{z}^{j}_{t},a^{j}_{t}\right)$
   else
     set $e^{i}_{t}=\emptyset$
   end if
   for iteration $1,2,3,\dots,K_{p}$ do
     Sample $N_{a}$ actions $a^{i}_{t:t+H}\sim\mathcal{N}\left(\mu^{i}_{t:t+H},{(\sigma^{i}_{t:t+H})}^{2}I\right)$
     Sample $N_{a}$ actions from actor $\hat{a}^{i}_{h}\sim\pi^{i,\textrm{Act}}(\cdot|o^{i}_{h},e^{i}_{h}),h=t:t+H$
     Get predictions by world model rollouts, $\Gamma=\{(\hat{z}^{i}_{h},a^{i}_{h},\hat{r}^{i}_{h})\}_{h=t:t+H}$
     Evaluate the trajectories by Eq. (7) and select top-$M$ elite action sequences
     Update action distribution following Eq. (8)
   end for
  end for
end for
Figure 12: Overview of implementation details. (Left): SEM Normalization. (Middle): Symlog transformation, where the $y$-axis represents the values encoded by two-hot encoding, and the $x$-axis represents the values aligned with the scalar target. (Right): Normalized $Q$-scale and return in the Dex-BottleCap task.
B.2 Model Architecture
The proposed sequential world model is composed of five components: encoder, dynamics predictor, reward predictor, critic, and actor, all implemented using MLPs. We report the network configurations and the number of parameters in each module on the Dex-BottleCap task $\left(|\mathcal{O}|=221,|\mathcal{A}|=26\right)$ in Table 2.
Table 2: The network configurations.
Module Dim Layers Act. Out Act. LayerNorm Param.
Encoder 512 2 Mish SEM Norm 189,952
Dynamics 512 2 Mish SEM Norm 553,984
Reward 512 2 Mish Linear 342,117
Critic 256 2 Mish Linear 736,970
Actor 256 2 Mish Tanh 210,996
We employ the Mish (Misra, 2020) activation function in the hidden layers, which is smooth and non-monotonic, ensuring stable gradient flow. For latent space construction, we adopt SEM Norm (Simplicial Embeddings Normalization) (Lavoie et al., 2022) to normalize the outputs of the encoder and dynamics predictor. It employs the softmax operator to project high-dimensional latent states onto a set of fixed-length simplices, thereby forming a soft sparse representation. This approach mitigates information degradation in high-dimensional latent spaces while preserving structural properties, striking a balance between expressive capacity and optimization feasibility. As shown in Figure 12 (left), the transformation of SEM Norm is defined as:
$\textrm{SEM}(z)=\oplus_{i=1}^{s}g^{(i)},\quad g^{(i)}=\frac{\exp(z_{i:i+L-1})}{\sum_{j=0}^{L-1}\exp(z_{i+j})},$ (11)
where $L$ is the dimension of the simplex, which is set to $L=8$ in our implementation, and $s=d_{z}/L$ is the number of segments of the simplex.
B.3 Reward and Value Prediction
Discrete Regression. The reward prediction and critic in SeqWM are modeled as discrete regression problems, which helps enhance robustness to extreme values. Take reward prediction for example, we discretize the reward range into $B$ intervals and use two-hot encoding (Bellemare et al., 2017; Hafner et al., 2025) to map the scalar reward to a $B$-dimensional vector. The two-hot encoding assigns weights to the two adjacent bins, reducing quantization errors and providing smoother gradients. Formally, for a scalar reward signal $r$, its two-hot encoding is defined as:
$\textrm{two-hot}(r)_{i}=\begin{cases}\nicefrac{{|b_{k+1}-r|}}{{|b_{k+1}-b_{k}|}},&\text{if }i=k\\ \nicefrac{{|b_{k}-r|}}{{|b_{k+1}-b_{k}|}},&\text{if }i=k+1\\ 0,&\text{else}\end{cases},k=\sum_{j=1}^{B}\mathbbm{1}(b_{j}\leq r).$ (12)
For training, we use the soft cross-entropy loss to optimize the predictor’s output:
$\textrm{Soft-CE}(\hat{r},r)=-\sum_{i=1}^{B}\textrm{two-hot}(r)_{i}\times\frac{\exp\left(\hat{r}_{i}\right)}{\sum_{j}\exp\left(\hat{r}_{j}\right)},$ (13)
Symlog and Symexp Transformations. We employ symmetric logarithmic (symlog) and exponential (symexp) transformations (Hafner et al., 2025) to align the predicted and target values. These transformations maintain continuity and differentiability while smoothing out large numerical values, allowing the model to suppress gradient instability caused by exceptionally large values, thereby enhancing numerical stability in high-variance scenarios. These transformations are defined as:
$\textrm{symlog}(x)=\textrm{sign}(x)\ln\left(|x|+1\right),\quad\textrm{symexp}(x)=\textrm{sign}(x)\left(\exp^{|x|}-1\right).$ (14)
Figure 12 (middle) illustrates how the symlog and symexp transformations enhance numerical stability. The symlog transformation is approximately linear near $x=0$ and logarithmic in the tails. This brings two stability benefits:
  1. 1)
    Sensitivity $\nicefrac{{dy}}{{dx}}=\nicefrac{{1}}{{1+|x|}}$ decreases as the target value increases, leading to nearly zero gradients for extremely large values.
  2. 2)
    Adaptive binning for two-hot targets. For a fixed width $\Delta x$, the corresponding original scale span is $\Delta y\approx(1+|x|)\Delta x$. Thus, extreme values are absorbed by the edge bins, preventing them from dominating the loss, while values near zero maintain fine granularity to preserve resolution.
Percentile Scaling of Critic
We employ a percentile-based scaling strategy to stabilize Q-value magnitudes. Specifically, we computes the dynamic range between the 5-th and 95-th percentiles of each batch, which effectively suppresses the influence of outliers. This range is updated smoothly using an EMA (Exponential Moving Average) controlled by a factor $\tau$:
$\delta^{(k+1)}=\tau\delta^{(k)}+(1-\tau)\cdot\left(p_{95}-p_{5}\right),$ (15)
where $p_{95}$ and $p_{5}$ are the 95-th and 5-th percentiles of the Q-values in the current batch, respectively, and $\tau$ is a smoothing factor (e.g., $\tau=0.99$).
Figure 12 (right) illustrates the empirical effect of percentile scaling on Q-value magnitudes. When the return rises rapidly at the beginning of training, the scaling coefficient also increases, adapting to the broader spread of Q-values while suppressing the destabilizing impact of extreme samples. As training progresses and the return gradually converges, the dynamic range of Q-values contracts and the scaling coefficient correspondingly decreases in a smooth manner, guided by the exponential moving average. This adaptive behavior ensures that the critic remains well-conditioned across different learning phases: it expands the effective range when exploration generates diverse values, and tightens the range when the policy stabilizes. Compared to naive normalization, this percentile-based strategy provides robustness against transient spikes or rare outliers, leading to more consistent value estimation and consequently more stable planning performance.
B.4 Hyperparameters
We summarize the hyperparameters used in SeqWM in Table 3 and Table 4.
Table 3: The Notations and Values of hyperparameters in the planner.
Hyperparameters Notations Value Hyperparameters Notations Value
rollout horizon $H$ 3 sampling actions $N_{p}$ 512
planning iterations $K$ 6 elites $M$ 64
temperature $\tau$ 0.5 actor samples $N_{a}$ 24
Table 4: The hyperparameters used in the world model.
Hyperparameters Value Hyperparameters Value Hyperparameters Value
batch size 1000 buffer size 1e6 dynamics coef 20
encoder lr scale 0.3 entropy coef 1e-4 lr 5e-4
n-step return 20 num bins 101 q coef 0.1
reward coef 0.1 step $\rho$ 0.5
Appendix C Additional Experiments
C.1 Scalability with the Number of Agents
Multi-Agent MuJoCo. To evaluate the scalability of SeqWM with respect to the number of agents, we conduct supplementary experiments on the 6a-Cheetah task in the MA-Mujoco (Multi-Agent MuJoCo) (Peng et al., 2021) environment. MA-Mujoco partitions a robot into multiple agents according to different body parts; in the 6a-Cheetah task, six agents are required to coordinate to control a single joint to run faster. To achieve scalability, we modify the communication protocol in Eq. (2) to transmit only the action sequence:
$e^{i+1}_{t}=e^{i}_{t}\oplus a^{i}_{t}$ (16)
The results, reported in Figure 14, show that SeqWM achieves an average return exceeding 12,000 on the 6-agent Cheetah task, which is, to our best knowledge, the state-of-the-art performance for this task.
Figure 13: Performance on 6a-Cheetah.
Figure 14: Success rates on Gate with more robots.
Gate with More Robots. We also evaluate scenarios with more robots on the Gate task, with success rates reported in Figure 14. As the number of robots increases, the task becomes more complex, but SeqWM still successfully learns effective cooperative strategies, achieving nearly 100% success rates in all settings within 4M time steps. Additionally, we visualize the learned behaviors of SeqWM with different numbers of robots in Figure 15, Figure 16, and Figure 17.
Figure 15: Visualization of the learned behaviors on Gate with 3 robots.
As shown in Figure 15, from t=1t{=}1 (Frame A) to t=2t{=}2 (Frame B) all agents head toward the passage with high $x$-velocity commands. Around $t{\approx}2$ (Frame B), priority is established: Robot 2 maintains forward $x$-command while Robot 1 and Robot 3 reduce theirs, preventing congestion (see “X Command”). From $t{=}2{\to}3$ (Frame B $\to$ C), Robot 1 clears the gate and issues a lateral $y$-command to move aside and create space (panel “Y Command”), with yaw staying near zero for all agents (“Robots’ Yaw”). Finally, from $t{=}3{\to}4$ (Frame C & D), Robot 1 passes as Robot 3 decelerates and holds position; Robot 3 proceeds last, completing a smooth, collision-free sequence (order: 2 $\rightarrow$ 1 $\rightarrow$ 3).
Figure 16: Visualization of the learned behaviors on Gate with 4 robots.
As shown in Figure 16, frames A–B show all agents accelerating toward the passage with high $x$-velocity commands. Around $t{\approx}3$ (Frame C), an implicit queue is established: Robot 2 keeps a near-unity $x$-command while Robot 1, Robot 3, and Robot 4 exhibit pronounced dips in the “X Command” panel, preventing blockage at the bottleneck. From $t{=}3{\to}4$ (Frame D), Robot 2 clears the gate first and issues a strong negative $y$-command to drift downward and vacate the corridor (“Y Command”), with yaw traces remaining close to zero for all robots (“Robots’ Yaw”). Robot 1 then increases its $x$-command and passes second while slightly steering upward in $y$ to avoid interference. Finally, from $t{=}4{\to}5$ (Frame E), Robot 3 proceeds third, followed by Robot 4, which had maintained the lowest $x$-command during the queuing phase, completing a smooth, collision-free sequence (order: 2 $\rightarrow$ 1 $\rightarrow$ 3 $\rightarrow$ 4).
Figure 17: Visualization of the learned behaviors on Gate with 5 robots.
In the 5-robot Gate task shown in Figure 17, the robots exhibit similar coordinated behaviors. Across agents, the X Command traces show pronounced wave-like undulations: troughs denote the robot waiting at the gate, whereas crests indicate the robot currently traversing the passage. Around $t{\approx}3$ (Frame C), an implicit queue emerges in the “X Command” panel: Robot 2 sustains a near-unity forward command while Robot 1 and Robot 3 reduce theirs moderately; Robot 4 and Robot 5 exhibit the deepest dips, preventing congestion at the bottleneck. From $t{=}3{\to}4$ (Frame D), Robot2 clears the gate first and issues a negative $y$-command to drift downward and vacate the corridor (“Y Command”), with yaw traces staying close to zero for all robots (“Robots’ Yaw”). Next, Robot 1 increases its $x$-command to pass second while slightly offsetting upward in $y$ to avoid interference. Finally, from $t{=}4{\to}5$ (Frame E), Robot 3 proceeds third, followed by Robot 4 and Robot 5, which had maintained the lowest $x$-commands during the queuing phase—completing a smooth, collision-free sequence (order: 2 $\rightarrow$ 1 $\rightarrow$ 3 $\rightarrow$ 4 $\rightarrow$ 5).
C.2 Additional Behavior Visualization
For further validation of the effectiveness of SeqWM, we provide additional visualizations of the learned behaviors on other tasks in Figure 18, Figure 19 and Figure 20.
Figure 18: Visualization of the learned behaviors on the Bi-DexHands tasks.
For BottleCap, SeqWM learns effective division of labor and cooperation, with one hand stably grasping the bottle body while the other rotates and successfully unscrews the cap without tilting or dropping the bottle. For CatchAbreast, both hands successfully catch the object in parallel positions. The hand responsible for catching starts moving to the right before the object falls, adjusting its position to make the catching process more stable. Overall, these visualizations indicate that SeqWM can learn stable policies in high-dimensional state and action spaces, achieving advanced cooperative behaviors that surpass baseline methods through sequential communication, enabling early prediction, role division, temporal alignment, and intent sharing.
Figure 19: Visualization of the learned behaviors on Gate task.
In Gate, from $t=1$ (Frame A) to $t=2$ (Frame B), both robots accelerate toward the narrow passage with high $x$-velocity commands. Around $t{\approx}3$ (Frame C), Robot 1 reduces its $x$-command while Robot 2 maintains forward motion, implicitly yielding priority to avoid collision. From t=3t{=}3 (Frame C) to t=4t{=}4 (Frame D), Robot 2 clears the gate and issues a negative $y$-command to move downward and create space (panel “Y Command”), while yaw remains near zero for both robots, indicating stable headings (panel “Robots’ Yaw”). Finally, from $t{=}4\to 5$ (Frame D $\to$ E), Robot 1 proceeds through the gate as Robot 2 decelerates, completing the task with smooth, collision-free coordination.
Figure 20: Visualization of the learned behaviors on Shepherd task.
For the Shepherd task, from $t=3$ (Frame C) to $t=4$ (Frame D), the sheep approaches the upper boundary. At this point, Robot 1 moves downward while Robot 2 advances to prevent the sheep from escaping, effectively guiding it back to the target area. This sequence highlights SeqWM’s predictive planning and adaptive role assignment: the leading robot influences the trajectory of the target robot, while the following robot adjusts accordingly to ensure successful shepherding.
C.3 Role of Actor and Planner
As shown in Eq. (2), the world model consists of both an actor and a critic, but their roles are different. The actor is only used to generate candidate actions for the planner and does not directly participate in decision-making as an explicit policy. In contrast, the critic is crucial because it estimates the value of candidate actions in Eq. (7) and guides the update of the action distribution in Eq. (8).
Table 5 presents the results of ablation studies that clearly demonstrate the complementary roles of the planner and actor. The planner-only setting, where the planner searches over randomly sampled actions without leveraging the estimates from the actor, achieves moderate performance across tasks but suffers from inefficiency and suboptimal exploration, as reflected in the consistent performance drop compared to the full planner-actor configuration. Conversely, the actor-only variant, which directly executes actions proposed by the actor without deliberation from the planner, performs even worse, indicating that the actor lacks the capability to independently handle task-level reasoning or long-term coordination. The planner-actor combination yields the best results in all tasks, showing that the actor contributes by narrowing the action search space toward promising regions, while the planner ensures robust long-term decision-making and correction of suboptimal actor proposals. This synergy allows the system to balance efficiency and accuracy, highlighting that the actor serves primarily as a prior for action generation, whereas the planner is responsible for structured search and task-level optimization.
Table 5: Ablation studies of actor.
Task Planner-Actor Planner-Only Actor-Only
BottleCap 480.4 $\pm$ 2.8 420.3 $\pm$ 15.8 372.5 $\pm$ 9.8
CatchAbreast 47.9 $\pm$ 0.8 43.0 $\pm$ 1.3 40.6 $\pm$ 1.0
Over2Underarm 30.5 $\pm$ 0.3 28.4 $\pm$ 0.9 27.5 $\pm$ 0.7
Over 50.1 $\pm$ 0.4 46.1 $\pm$ 1.5 30.5 $\pm$ 1.1
Pen 221.5 $\pm$ 2.2 190.1 $\pm$ 6.6 163.8 $\pm$ 5.0
Scissors 686.3 $\pm$ 3.1 634.6 $\pm$ 22.1 592.1 $\pm$ 17.1
C.4 Early-Stopping Planner
Inference Time Cost. We report the per-step execution time of SeqWM on the BottleCap task using a single RTX A6000 GPU on the left side of Figure 21. The execution time increases almost linearly with the rollout horizon $H$ and the number of planner iterations $K$, which is consistent with the design of SeqWM. With the default settings, SeqWM achieves a per-step execution time of 12.8 ms, making it suitable for most real-time robotic tasks.
Early-Stopping Heuristic. To further enhance the efficiency of SeqWM, we introduce an early-stopping heuristic in Section 4.2 that terminates iterations when the change in the action distribution is not significant. The KL divergence is used as a measure of distribution change, and the execution time and performance under different thresholds on the BottleCap task are shown on the right side of Figure 21. When the threshold is set to 0.5, SeqWM reduces the execution time by approximately 27.3% while incurring only about 5.9% performance loss.
Figure 21: The per-step execution time of SeqWM. (Left): Time cost under different rollout horizons $H$ and planner iterations $K$. (Right): Time cost and performance with and without early-stopping heuristic.
C.5 Sim-to-Real Deployment
We implement all three Multi-Quad tasks in an $8\mathrm{m}\times 5\mathrm{m}$ indoor space. Each task involves two Unitree Go2-W quadruped robots. The room is equipped with eight Mars cameras, and real-time localization of robots and objects is provided by the NOKOV 3D motion capture system.
  • PushBox: We use a cardboard box of $1.2\mathrm{m}\times 1.2\mathrm{m}\times 0.5\mathrm{m}$ and approximately $6~\mathrm{kg}$ in mass. The box is sufficiently large that a single robot cannot independently control its movement direction, making cooperation essential. The static and kinetic friction coefficients between the box and the ground are both approximately $0.5$.
  • Gate: A $1\mathrm{m}$-wide doorway is set up. As shown in Figure 7 (b)-A, the two robots cannot pass through side-by-side, requiring coordinated navigation.
  • Shepherd: A DJI EP robot acts as the guided agent (sheep). It is equipped with an omnidirectional chassis to simulate sheep behavior: it moves away from the nearest herding robot and its speed is inversely proportional to the distance to that robot.
Figure 22: Real-world setups.
We employ the following strategies to enhance the generalization capability of SeqWM and facilitate sim-to-real transfer:
  • Observation transformation: Positions of other robots are transformed from the global frame into the ego-centric frame of the current robot, reducing observation complexity and improving policy generalization.
  • Domain randomization: Taking the PushBox task as an example, we randomize the initial positions/orientations of both robots and the box, the position and distance of the target, and the friction coefficient between the box and floor to improve robustness to environmental variations.
  • Sensor and actuation perturbations: Random noise is added to sensor readings, and small delays with noise are introduced into control commands to emulate real-world sensing errors and actuation inaccuracies.